### v2

# MOLLM: Multi-Objective Large Language Model for Molecular Design – Optimizing with Experts

Abstract
Molecular design plays a critical role in advancing fields such as drug discovery, materials science, and chemical engineering. This work introduces the Multi-Objective Large Language Model for Molecular Design (MOLLM), a novel framework that combines domain-specific knowledge with the adaptability of large language models to optimize molecular properties across multiple objectives. Leveraging in-context learning and multi-objective optimization, MOLLM achieves superior performance and innovation, consistently surpassing state-of-the-art (SOTA) methods. We significantly improve the efficiency of our framework, making it 14 times faster and substantially more cost-effective without compromising performance compared to the latest similar work. Our results demonstrate that MOLLM consistently outperforms SOTA models across experiments and excels on the PMO benchmark. In addition, we provide extensive ablation studies and analysis to evaluate the effectiveness of each component and the quality of the output molecules.
1 Introduction
Molecular design is fundamental in fields such as drug discovery, materials science, and chemical engineering. In these areas, the ability to design novel molecules with targeted properties, including stability, reactivity, or bioactivity, can drive significant advancements, from the development of new pharmaceuticals to the creation of sustainable, innovative materials. Traditionally, molecular design has relied on trial-and-error experimentation and repeated synthesis, which is resource intensive, time-consuming, and ultimately inefficient. During the past few decades, with rapid advances in computational power, various machine learning techniques Elton et al. (2019); Du et al. (2022) have been introduced not only to accelerate this process but also to enable the discovery of novel, more potent molecules. Methods include Bayesian Optimization (BO) Tripp et al. (2021), Multi-Objective Optimization (MOO) Liu et al. ; Choi et al. (2023); Verhellen (2022), Markov Chain Monte Carlo (MCMC) Xie et al. (2021); Sun et al. (2022), Genetic Algorithms (GA) Jensen (2019); Nigam et al. (2019); Liu et al. (2024b); Brahmachary et al. (2024); Wang et al. (2024b, a), Reinforcement Learning (RL) Olivecrona et al. (2017); Jin et al. (2020); Fu et al. (2022), and Deep Learning (DL) models Jin et al. (2018b); Bagal et al. (2021); Lee et al. (2023); Fang et al. (2024).
Although these methods have yielded excellent results, most of them lack the integration of expert knowledge during runtime, despite the crucial role of professional feedback and search direction in molecular design. Large Language Models (LLMs), typically based on transformer architectures Vaswani (2017), are pre-trained on extensive high-quality data, including books and academic papers, enabling them to capture domain-specific expertise. They have demonstrated significant potential in scientific discovery, particularly in molecular understanding and the generation of novel molecular candidates, as exemplified by models like GPT-4 AI4Science and Quantum (2023). Recent studies highlight the advantages of in-context learning Nguyen and Grover (2024) and iterative evolutionary approaches Wang et al. (2024b) in enhancing LLM effectiveness. However, research in this area remains nascent, with only preliminary findings and a lack of systematic investigation.
Furthermore, despite significant progress in training large neural networks to understand chemistry and molecular structures with domain knowledge, these models often require additional parameters and retraining, particularly for MOO, as seen in MolGPT Bagal et al. (2021) and LICO Nguyen and Grover (2024). In contrast, MOLLEO Wang et al. (2024b) leverages domain knowledge from pre-trained large language models without additional training but still relies on GB-GA within its framework and only uses a weighted sum for solving multi-objective tasks. In practice, most molecular design tasks optimizes multiple objectives, yet existing methods often ignore this aspect. For example, GB-BO Tripp et al. (2021), JTVAE Jin et al. (2018a), and MolGen Fang et al. (2024) provide limited multi-objective capabilities. MolGPT requires specific training for different objectives, restricting its flexibility, while MolGen focuses primarily on target molecular discovery and employs only single-objective optimization.
To address these gaps and enhance multi-objective molecular design, we propose Multi-Objective Large Language Model (MOLLM), a LLM-based framework that integrates MOO, in-context learning and prompt engineering. Our model is mainly consisted of a mating module to generate parent molecules for in-context learning, a prompt template to integrates all information and instructions to maximally leverage the knowledge in LLM, a experience pool, and a selection module that contains both Pareto front selection and fitness value selection. The results show that our model demonstrates SOTA performance on different objectives, especially in multi-objective cases and when the number of objectives become larger. Our key contributions are:
We carefully design the in-context learning and prompt engineering mechanisms to fully leverage the domain knowledge encoded in LLMs. These components are seamlessly integrated into the MOO framework, achieving SOTA performance on the PMO benchmark and additional multi-objective optimization tasks in terms of both quality and efficiency. Our framework requires no task-specific training, benefits from domain reasoning capabilities, and is adaptable to various LLMs. Unlike MOLLEO, our model employs LLMs for all mating operations, making the entire framework LLM-driven. This model is easy to use and can effectively accelerate the process of discovering novel drugs.
We develop an experience pool that is refined iteratively during optimization. This memory mechanism accelerates convergence and improves optimization results by enhancing the model’s exploitation ability. Moreover, it improves the explainability of the optimization process and provides reusable and insightful experience at the end of each experiment.
We introduce several modifications to enhance exploration and balance it with exploitation: (1) We identify that the number of output molecules clearly influences the LLM’s exploratory capacity and determine an optimal output number (illustrated in Section 7.3 in Appendix). (2) We employ Pareto front selection instead of only using simple objective summation, which promotes diversity in search directions. These components contribute to increased performance as the number of objectives grows compared to MOLLEO. By further reducing redundant components in the framework, we achieve approximately 14× faster than MOLLEO and requires significantly fewer LLM queries, without compromising performance.
2 Related Work
2.1 Molecular Design with Machine Learning
Numerous advanced models for molecular design span GA, BO, MOO, MCMC, RL, and DL methodologies. Deep Learning (DL) leverages neural networks in various molecular design models. Differentiable Scaffolding Tree (DST) Fu et al. (2021) with GNNs, Junction Tree Variational Autoencoders (JTVAE) Jin et al. (2018a), and VJTNN+GAN Jin et al. (2018b) combine generative and adversarial architectures to generate molecules. MOOD Lee et al. (2023) utilizes Diffusion models to address out-of-distribution generation. Recent developments in Generative Pre-trained Transformers (GPT) led Bagal et al. to train MolGPT Bagal et al. (2021) on next-token prediction, while Fang et al. pre-trained MOLGEN Fang et al. (2024) on molecule reconstruction tasks, achieving cross-domain applicability. Although DL methods offer powerful capabilities in capturing complex molecular structures and enabling cross-domain applicability such as DST, JTVAE, and MolGPT, they often underperform in MOO scenario. Latent Space Optimization (LSO) Abeer et al. (2024) has further advanced multi-objective molecular design, but only for deep generative models.
Reinforcement Learning (RL) combined with DL iteratively refines molecules by learning from feedback, often based on property scores. REINVENT Olivecrona et al. (2017) applies RL to train an RNN to generate molecules meeting multiple goals, while RationaleRL Jin et al. (2020) uses a Graph Neural Network (GNN) to generate molecules by building interpretable substructures, or “rationales”. Based on REINVENT, Shin et al. proposed a novel divide-and-conquer approach called DyMol Shin et al. (2024) to train the model for multiple objectives and achieve SOTA results. Kim et al. also achieve SOTA performance by integrating genetic algorithms into GFlowNets Kim et al. (2024). Guo and Schwaller achieve SOTA performance by introducing Augmented Memory, which integrates experience replay and SMILES augmentation for efficient molecular optimization Guo and Schwaller (2024).
In addition to DP methods, classical probabilistic models and optimization methods also achieve SOTA performance in many cases, such as in PMO Gao et al. (2022). A notable example of Genetic Algorithms (GA) is GB-GA Jensen (2019), commonly used as a baseline, where molecular structures are modified in graph form during mating operation. AkshatKumar et al. Nigam et al. (2019) introduced a neural network discriminator to enhance diversity, surpassing GB-GA in maximizing penalized-logP Gómez-Bombarelli et al. (2018). Later, Tripp et al. Tripp et al. (2021) employed a Tanimoto kernel in a Gaussian Process in GB-GA, outperforming GB-GA. It uses SELFIES Krenn et al. (2020), a 100% valid molecular representation system; however, Gao et al. Gao et al. (2022) later showed there are no obvious shortcomings of SMILES compared to SELFIES. MLPS Liu et al. combines MOO with BO and an encoder-decoder network to efficiently locate global Pareto-optimal solutions, while Verhellen et al. introduced a graph-based MOO Verhellen (2022) for molecular optimization. Furthermore, MARS Xie et al. (2021) uses Markov Chain Monte Carlo (MCMC) to explore chemical spaces probabilistically to identify molecules with desirable properties. Similarly, MolSearch Sun et al. (2022) utilizes Monte Carlo tree search for multi-objective molecular discovery. However, GA, BO, MOO, and MCMC methods are independent of domain knowledge, which is highly beneficial in molecular design but challenging to incorporate into such algorithms.
2.2 Multi-Objective Optimization and Genetic Algorithm with LLM
Recently, Large Language Models (LLMs) have demonstrated remarkable performance across various Natural Language Processing (NLP) benchmarks Brown (2020); AI4Science and Quantum (2023), sparking interest in their application as optimization operators in MOO to address the challenges of high-dimensional search spaces and to incorporate domain knowledge Wu et al. (2024a). For instance, OPRO Yang et al. (2024) and LMEA Liu et al. (2024b) employ LLMs as crossover and mutation operators within GA, using prompts that include parent values from the current population, with LMEA further exploring the balance of exploitation and exploration by adjusting the temperature parameter. Furthermore, Wang et al. Wang et al. (2024c) investigated constrained MOO with prompt engineering, demonstrating promising alignment results. Other studies have highlighted the effectiveness and efficiency of LLMs in GA compared to standalone LLMs and traditional MOO algorithms, proposing well-structured pipelines Liu et al. (2023, 2024a, 2024c); Huang et al. (2024); Brahmachary et al. (2024). However, research on LLMs with MOO is still nascent, with results largely limited to preliminary findings in numerical optimizations and planning problems.
2.3 Molecular Design with LLM
LLMs with pre-trained domain knowledge are increasingly popular for accelerating drug discovery and materials design AI4Science and Quantum (2023). In particular, ChemCrow M. Bran et al. (2024) uses LLMs as agents capable of reasoning, planning, and selecting appropriate external tools to iteratively refine candidates in chemical tasks. LICO Nguyen and Grover (2024) improves molecule generation through in-context learning by pretraining the model with separate embedding and prediction layers, while Moayedpour et al. Moayedpour et al. (2024) extend this approach to multi-objective setups, and MolReGPT Li et al. (2024) advances few-shot learning for molecular optimization. MOLLEO Wang et al. (2024b) applies GA combined with LLMs for molecular design, aligning with our framework, but differing significantly in details. MOLLEO’s results as well as research in prompts remain preliminary, lacking specific consideration and modification for multi-objective settings. Prompt-MolOpt Wu et al. (2024b) explicitly leverages the domain knowledge of LLMs for molecular design and demonstrates strong zero- and few-shot capabilities through property-specific prompt embeddings combined with a sequence-to-sequence Transformer trained on substructure-annotated molecule pairs. However, its practical evaluations are mostly limited to two-objective settings, and its performance on more complex combinations still depends on property-specific prompt engineering and single-objective training data.
3 Methodology
The core ideas of MOLLM is that molecular design should leverage prior domain knowledge embedded in SOTA LLMs rather than training models from scratch, disregarding expert feedback during optimization, or relying on external algorithms such as GB-GA as operators. Therefore, we propose utilizing LLMs exclusively for both crossover and mutation operations in our model. The reason of two operations is to balance exploitation and exploration. While LLMs may not always generate molecules that perfectly consider the trade-offs of objectives, we incorporate Pareto front selection within (MOO) to ensure that molecules selected for the next generation better account for all objectives while maintaining structural diversity. Empirical experiments demonstrate that well-formulated prompts and in-context learning significantly enhance the utilization of LLM knowledge and the information encoded in parent molecules. Thus, we carefully design a prompt template comprising five key components, making it adaptable to any LLM.
3.1 MOLLM Overview
Figure 4 presents the complete MOLLM optimization pipeline. The task involves unconstrained molecular optimization, where, given a set of objectives, the model is initialized with molecules selected from the ZINC250K dataset. It then iterates through mating, prompt generation, scoring, experience updating, and next-generation selection. Initialization: For an optimization problem with one or multiple objectives, we initialize with molecules, either randomly selected or chosen based on the best or worst objective values from the ZINC250K dataset. In our setting, . The ZINC dataset Irwin et al. (2012) is selected as it is widely used for population initialization in molecular optimization studies Wang et al. (2024b); Jensen (2019); Fu et al. (2022). ZINC250K comprises approximately 250,000 curated drug-like molecules from the ZINC database, providing key properties such as chemical structures, logP, QED, and SA, making it well-suited for drug discovery and molecular optimization tasks.
Mating: This step involves prompting the LLM to generate new candidate molecules that are expected to improve on the parent molecules given. First, the parent molecules are randomly selected from the current population with probabilities and for crossover and mutation, respectively. Each crossover involves two parents, while mutation involves a single parent. These selected parents are then formatted into a flexible prompt template. Our prompt template consists of five key components: multi-objective requirements, objective descriptions, parent objective values (for in-context learning), output instructions, and past experience. The whole framework of our model and an example is shown in Figure 4. In this setup, the model receives structured input specifying primary objectives, descriptions of molecular modifications (in SMILES format) that may increase or decrease property values, and parent molecules represented by their SMILES structures, objective values, and an aggregated objective score as an overall performance indicator. The output instructions specify that only molecular structures should be generated, omitting explanations to significantly reduce runtime and query costs without affecting performance. We employ crossover and mutation to balance exploitation and exploration. Although LLMs perform well in crossover due to their straightforward nature, they struggle with mutation, as its prompt is highly similar to crossover. To address this, we provide a list of common molecular mutation operations in the instructions to improve exploration. Finally, after generating offspring, we identify the best and worst molecules among them and query the LLM to update its experience based on these molecules and experience in the last iteration. This iterative refinement allows the experience to evolve dynamically, transitioning from general suggestions to more detailed and actionable guidance over time.
Multi-objective optimization: At this stage, we typically have parent molecules from the previous generation and offspring from the current generation, assuming all molecules are valid, where denotes the population size. These molecules are then combined and subjected to either Pareto front selection or F-value selection, where F represents the sum of normalized objective values, to determine the top candidates for the next generation. The selection operation is executed with equal probability (50% each) for both methods. This hybrid selection strategy balances exploration and exploitation: F-value selection allows the model to focus on the current optimal solutions, while Pareto front selection promotes diversity in the next generation, reducing the risk of premature convergence to a local optimum.
4 Experiment
4.1 Task settings
The formulation of MOO using GA has often lacked rigor in previous studies. First, oracle calls should be restricted to ensure fair comparisons and practical applicability, since the evaluation of certain molecular properties requires costly experiments or specifically trained models, as noted in the Practical Molecular Optimization (PMO) benchmark Gao et al. (2022). Additionally, the initial population significantly impacts final performance under a fixed number of oracle calls, yet this factor has been largely overlooked in methods that incorporate genetic algorithms, such as MARS, MOLLEO, and GB-GA. In practical applications, researchers often initialize the search with the best available molecules rather than selecting them entirely at random. To ensure a comprehensive and fair evaluation of our model, we conduct experiments on three distinct initialization scenarios: the top 100, bottom 100, and randomly sampled molecules from the ZINC 250K dataset, using their F-values as indicators. The best-initialization scenario assesses the model’s upper performance limit, the random initialization reflects common real-world usage, and the worst-initialization scenario presents a more challenging optimization task. We adhere to the PMO benchmark and operate within a budget of 5,000 oracle calls. For molecular property optimization, we adopt the multi-objective task from MOLLEO: maximizing QED (drug-likeness) and JNK3 (c-Jun N-terminal Kinase 3), and minimizing SA (synthetic accessibility), LogP (octanol-water partition coefficient), DRD2 (dopamine receptor D2 affinity), LogS (log of solubility), reduction potential, and GSK3 (Glycogen Synthase Kinase 3 Beta). In addition to these well-defined objectives, we also include maximizing BBBP (Blood-Brain Barrier Permeability), a more complex and less predictable property influenced by multiple biological factors. In addition, we evaluated our method on the PMO benchmark, which includes structure-based optimization, name-based optimization, and property optimization. This benchmark features both single-objective tasks and multi-objective scenarios, such as the multi-parameter optimization of amlodipine, modeled after Amlodipine, a calcium channel blocker. This example task focuses on optimizing QED, logP, HBA/HBD, and SA, aiming to generate compounds that are realistic drug candidates while structurally distinct from the original molecule.
4.2 Metrics
To fully evaluate the performance in many aspects, we use several metrics. The most important goal is maximizing the sum of normalized property values, denoted as F value, representing the absolute improvement that accountsnts for all the objectives. On top of that, we use uniqueness, validity, diversity and efficiency to full evaluate the ability of model to propose molecules. However, these additional metrics need to be considered in conjunction with the F-value, as it is less meaningful of other metrics if they have relatively low F values.
Top 1 F & Mean Top 10 F: F (fitness) is the sum of the normalized objective values, which gives the direct representation of the strength of a molecule Wang et al. (2024b). The weight in our experiment to each objective is the same.
where is a molecule in SMILES form, k is the number of objectives, and is the weight and normalized objective value. If an objective is to be minimized, it will be transformed by . We give an equal weight to each objective. Uniqueness: the fraction of valid generated molecules that are unique. A low uniqueness highlights repetitive molecule generation and a low level of distribution learning by the model Bagal et al. (2021), while a high uniqueness value means that the model effectively explores novel molecules. Validity: the fraction of molecules generated that are valid, it measures how well the model has learned the SMILES grammar and the valency of atoms Bagal et al. (2021). Structural Diversity: Structural diversity reflects the chemical diversity of the Pareto set and is computed by taking the average pairwise Tanimoto distance between Morgan fingerprints of molecules in the set Benhenda (2017). Efficiency: Efficiency is compared by the running time in hours, as well as LLM calls if application. It is a important metric when using LLM for inference, because querying LLM incurs high computational costs.
4.3 Baselines
To demonstrate the superiority and for fair comparison extensively, we choose SOTA models from a series of algorithms including GA, BO, MCMC, RL, DL and LLM-based method as our baselines. These algorithms are GB-GA, GB-BO, JT-VAE, MARS, REINVENT, MOLLEO, and recently proposed DyMol and Genetic-GFN which have achieved SOTA performance. For a fair comparison, we use Chatgpt 4o for both MOLLM and MOLLEO. We use the default hyperparameters for GB-GA, JT-VAE, GB-BO, MARS, REINVENT defined in PMO benchmark Gao et al. (2022). In terms of MOLLEO, DyMol and Genetic-GFN, we also use the default hyperparameters defined in their codes and papers. For fair comparison, the normalized objectives are applied for all methods, which also includes the correct optimizing direction. Each model is run for 5 seeds independently and takes the average. On the PMO benchmark, we use the scores evaluated in the MOLLEO Wang et al. (2024b) paper, including the Augmented memory. We also notice the GuacaMol Brown et al. (2019) benchmark, but it does not constrain the oracle calls, which is unsuitable for our experiments.
4.4 Main Experiment Results
Metric | GB-GA | JT-VAE | GB-BO | MARS | REINVENT | MOLLEO | DyMol | Genetic-GFN | MOLLM(ours) |
(Worst initial) | |||||||||
Top1 F | 4.048 | 3.838 | 3.665 | 3.891 | - | 4.096 | - | - | 4.229 |
Top10 F | 4.019 | 3.784 | 3.647 | 3.852 | - | 4.044 | - | - | 4.186 |
AUC-Top10 | 3.789 | 3.712 | 3.489 | 3.740 | - | 3.825 | - | - | 3.915 |
Hypervolume | 0.474 | 0.364 | 0.167 | 0.488 | - | 0.720 | - | - | 0.737 |
Uniqueness | 0.786 | 1.000 | 1.000 | 0.488 | - | 0.672 | - | - | 0.829 |
Validity | 1.000 | 1.000 | 1.000 | 1.000 | - | 0.930 | - | - | 0.937 |
Diversity | 0.583 | 0.847 | 1.000 | 0.826 | - | 0.656 | - | - | 0.603 |
(Random initial) | |||||||||
Top1 F | 4.017 | 3.874 | 4.003 | 3.931 | 4.230 | 4.190 | 4.232 | 4.299 | 4.413 |
Top10 F | 3.975 | 3.831 | 3.968 | 3.868 | 4.136 | 4.076 | 4.164 | 4.262 | 4.388 |
AUC-Top10 | 3.861 | 3.771 | 3.802 | 3.789 | 3.930 | 3.949 | 4.001 | 4.098 | 4.116 |
Hypervolume | 0.643 | 0.428 | 0.507 | 0.409 | 0.742 | 0.860 | 0.868 | 0.871 | 0.905 |
Uniqueness | 0.821 | 0.956 | 1.000 | 0.477 | 0.690 | 0.575 | 0.986 | 0.349 | 0.872 |
Validity | 1.000 | 1.000 | 1.000 | 0.999 | 0.979 | 0.938 | 1.000 | 0.998 | 0.908 |
Diversity | 0.623 | 0.778 | 0.717 | 0.819 | 0.640 | 0.570 | 0.581 | 0.633 | 0.441 |
(Best initial) | |||||||||
Top1 F | 4.583 | 4.329 | 4.605 | 4.419 | - | 4.699 | - | - | 4.699 |
Top10 F | 4.582 | 4.132 | 4.467 | 4.181 | - | 4.564 | - | - | 4.628 |
AUC-Top10 | 4.130 | 4.091 | 4.237 | 4.137 | - | 4.362 | - | - | 4.481 |
Hypervolume | 0.968 | 0.917 | 1.275 | 0.975 | - | 1.168 | - | - | 1.175 |
Uniqueness | 0.729 | 1.000 | 1.000 | 0.432 | - | 0.678 | - | - | 0.942 |
Validity | 1.000 | 1.000 | 1.000 | 0.999 | - | 0.913 | - | - | 0.790 |
Diversity | 0.424 | 0.792 | 0.630 | 0.788 | - | 0.600 | - | - | 0.491 |
1 Objective | 2 Objectives | 3 Objectives | 4 Objectives | 5 Objectives | 6 Objectives | |
Metric | MOLLM MOLLEO | MOLLM MOLLEO | MOLLM MOLLEO | MOLLM MOLLEO | MOLLM MOLLEO | MOLLM MOLLEO |
Top1 F | 0.948 0.941 | 1.901 1.887 | 2.901 2.891 | 3.901 3.890 | 4.413 4.098 | 5.183 4.964 |
Top10 F | 0.948 0.936 | 1.901 1.882 | 2.901 2.886 | 3.901 3.887 | 4.388 4.065 | 5.164 4.948 |
Uniqueness | 0.929 0.150 | 0.666 0.231 | 0.778 0.273 | 0.807 0.387 | 0.872 0.575 | 0.957 0.591 |
Validity | 0.796 0.159 | 0.962 0.552 | 0.946 0.803 | 0.946 0.783 | 0.908 0.938 | 0.890 0.926 |
Diversity | 0.538 0.865 | 0.450 0.646 | 0.510 0.627 | 0.375 0.614 | 0.441 0.573 | 0.529 0.611 |
Following the experimental settings of MOLLEO Wang et al. (2024b), we first conduct experiments to optimize five molecular properties simultaneously using molecules sampled from the ZINC 250K dataset. Among these objectives, three are minimized: SA, DRD2, and GSK3, while two are maximized: QED and JNK3. Each model is run with five different random seeds, and the final results are reported as the average over these runs. Since the initial population for REINVENT, DyMol, and Genetic-GFN cannot be explicitly set, these models are only evaluated in the randomly initialized scenario. The key evaluation metrics are top-1 fitness and average top-10 fitness, both of which directly reflect the sum of the normalized property values. To enhance clarity, the highest values in each metric are highlighted in Table 1. Our model demonstrates a significant improvement over other SOTA models across all three initialization cases, with a clear performance gap compared to the second-best approach. In both random-initialization and worst-initialization scenarios, even the mean top-10 F-value of MOLLM exceeds the top-1 F-value of the second-best model, highlighting the superior performance and convergence capabilities of our approach. We evaluate the novelty of our output molecules to ZINC 250K dataset. Notably, every run for all scenarios outputs 100% novel molecules that do not exist in ZINC 250K.
In the best-initialization scenario, while the top-1 fitness of MOLLEO matches that of MOLLM, the mean top-10 fitness of MOLLM is noticeably higher than both the top-1 and mean top-10 fitness of all other models. Furthermore, our model maintains a uniqueness rate above 90%, whereas MOLLEO, despite being another LLM-based method, exhibits significantly lower uniqueness. This underscores the strong capability of MOLLM in effectively exploring the chemical space. The validity of generated molecules is also comparable to other models. Although our model exhibits relatively lower diversity among the top-100 molecules, we observe that models with higher diversity often achieve lower top fitness values. This suggests that direct comparisons of diversity is less meaningful in this context but highlight a potential direction for future improvements.
4.5 Property values of 5-objective optimization
Method | SA | DRD2 | GSK3 | QED | JNK3 |
Starting population | |||||
(random sampled) | 3.021 | 0.009 | 0.026 | 0.730 | 0.015 |
MOLLM | 1.990 | 0.001 | 0.069 | 0.828 | 0.646 |
MOLLEO | 1.861 | 0.001 | 0.013 | 0.919 | 0.133 |
Genetic-GFN | 1.998 | 0.001 | 0.018 | 0.865 | 0.408 |
DyMol | 2.125 | 0.001 | 0.019 | 0.763 | 0.409 |
REINVENT | 1.878 | 0.001 | 0.012 | 0.900 | 0.235 |
MARS | 2.046 | 0.002 | 0.010 | 0.862 | 0.010 |
GB-BO | 1.725 | 0.001 | 0.018 | 0.887 | 0.111 |
JT-VAE | 2.144 | 0.001 | 0.008 | 0.904 | 0.018 |
GB-GA | 1.670 | 0.001 | 0.036 | 0.901 | 0.107 |
We present the objective values for the five-objective optimization task in Table 10. Compared to the initial population by random sampling, all models achieve notable improvements across nearly all objectives. Notably, while MOLLM achieves competitive performance on SA, DRD2, GSK3, and QED, it demonstrates a significant advantage in optimizing JNK3, outperforming the second-best model by a substantial margin. This result highlights MOLLM’s ability to effectively balance conflicting objectives, as improving JNK3 alone is relatively straightforward, but simultaneously enhancing JNK3 while reducing DRD2 and GSK3 remains a much more challenging task, which the other models fail to address adequately.
4.6 Results on PMO
Task type | Objective() | REINVENT | Augmented Memory | Graph GA | GP BO | MOLLEO | Genetic GFN | MOLLM (Ours) |
Property optimization | QED | 0.941 | 0.941 | 0.940 | 0.937 | 0.948 | 0.942 | 0.943 |
JNK3 | 0.783 | 0.773 | 0.553 | 0.564 | 0.790 | 0.764 | 0.822 | |
DRD2 | 0.945 | 0.962 | 0.964 | 0.923 | 0.968 | 0.974 | 0.974 | |
GSK3 | 0.865 | 0.889 | 0.788 | 0.851 | 0.863 | 0.881 | 0.836 | |
Name-based optimization | mestranol_similarity | 0.618 | 0.764 | 0.579 | 0.627 | 0.972 | 0.708 | 0.904 |
albuterol_similarity | 0.896 | 0.918 | 0.874 | 0.902 | 0.985 | 0.949 | 0.985 | |
thiothixene_rediscovery | 0.534 | 0.562 | 0.479 | 0.559 | 0.727 | 0.583 | 0.547 | |
celecoxib_rediscovery | 0.716 | 0.784 | 0.582 | 0.728 | 0.864 | 0.802 | 0.869 | |
troglitazone_rediscovery | 0.452 | 0.556 | 0.377 | 0.405 | 0.562 | 0.511 | 0.707 | |
perindopril_mpo | 0.537 | 0.598 | 0.538 | 0.493 | 0.600 | 0.595 | 0.646 | |
ranolazine_mpo | 0.760 | 0.802 | 0.728 | 0.735 | 0.769 | 0.819 | 0.843 | |
sitagliptin_mpo | 0.021 | 0.479 | 0.433 | 0.186 | 0.584 | 0.634 | 0.587 | |
amlodipine_mpo | 0.642 | 0.686 | 0.625 | 0.552 | 0.773 | 0.761 | 0.704 | |
fexofenadine_mpo | 0.769 | 0.686 | 0.779 | 0.745 | 0.847 | 0.856 | 0.856 | |
osimertinib_mpo | 0.834 | 0.856 | 0.808 | 0.762 | 0.835 | 0.860 | 0.888 | |
zaleplon_mpo | 0.347 | 0.438 | 0.456 | 0.272 | 0.510 | 0.552 | 0.571 | |
median1 | 0.372 | 0.335 | 0.287 | 0.325 | 0.352 | 0.379 | 0.339 | |
median2 | 0.294 | 0.290 | 0.229 | 0.308 | 0.275 | 0.294 | 0.310 | |
Structure-based optimization | isomers_c7h8n2o2 | 0.842 | 0.954 | 0.949 | 0.662 | 0.984 | 0.969 | 0.977 |
isomers_c9h10n2o2pf2cl | 0.642 | 0.830 | 0.719 | 0.469 | 0.874 | 0.897 | 0.893 | |
deco_hop | 0.666 | 0.688 | 0.619 | 0.629 | 0.942 | 0.733 | 0.938 | |
scaffold_hop | 0.560 | 0.565 | 0.517 | 0.548 | 0.971 | 0.615 | 0.967 | |
valsartan_smarts | 0.000 | 0.000 | 0.000 | 0.000 | 0.867 | 0.135 | 0.904 | |
Total () | 14.036 | 15.356 | 13.823 | 13.182 | 17.862 | 16.213 | 18.010 | |
Rank () | 5 | 4 | 6 | 7 | 2 | 3 | 1 |
On the PMO benchmark, we comprehensively evaluate our model across both single-objective and multi-objective optimization tasks, including property, name-based, and structure-based optimization, as shown in Table 4. The best values are highlighted in bold, with the next top three values underlined. Our model ranks first in 12 out of 23 tasks and achieves the highest total score and overall ranking, demonstrating its superiority across diverse objectives and consistent SOTA performance.
5 Empirical Study
In addition to the SOTA results from our main experiments involving the optimization of five objectives, we conduct further experiments with one to six objectives to assess the efficacy of MOLLM across varying optimization complexities and less predictable properties. Following this, we present an analysis of an interesting finding related to the experience pool utilized in our algorithm. We also conducted experiments of the effects of different output molecules per query in Section 7.3 in the Appendix.
5.1 More Objectives
To further investigate MOLLM’s performance across different objective configurations, we conduct experiments using random initialization across scenarios with one to six objectives, starting with only QED to QED + SA + DRD2 + GSK3 + JNK3 + BBBP. As the number of objectives increases, the performance gap between MOLLM and MOLLEO widens, particularly when optimizing more than four objectives, highlighting the superior capability of MOLLM in handling MOO. Additionally, MOLLM consistently achieves higher uniqueness and competitive validity compared to MOLLEO, while in MOLLEO these metrics tend to degrade significantly when optimizing fewer objectives. The consistently high uniqueness across all cases underscores the stability and effectiveness of MOLLM in optimization tasks with varying numbers of objectives. To further assess the robustness of MOLLM, we introduce BBBP (Blood-Brain Barrier Permeability) as a sixth objective, as it is a more complex and less predictable property with limited domain knowledge. Notably, despite the increased difficulty, MOLLM successfully generates a top 100 molecule set where all molecules are Blood-Brain Barrier Permeable (BBB+), demonstrating its strong adaptability and effectiveness in optimizing challenging molecular properties.
Method | MOLLM | MOLLEO | Graph GA | Gp-BO | Genetic-GFN | MARS | JT-VAE | DyMol | REINVENT |
Running time(hours) | 0.52 | 7.32 | 1.04 | 0.68 | 0.45 | 0.69 | 3.08 | 0.28 | 0.52 |
LLM calls | 2908 | 8517 | - | - | - | - | - | - | - |
Apart from that, without early stopping, MOLLM only uses nearly LLM calls compared to MOLLEO, more than even 14x faster than MOLLEO in run time to achieve significantly better results, as shown in Table 5. The running time is also competitive to other methods that are not based on LLMs.
5.2 Experience Pool
Top1 F | Top10 F | AUC-Top10 | Hypervolume | Uniqueness | Validity | Diversity | |
0.0 | 4.276 | 4.245 | 3.992 | 0.867 | 0.949 | 0.900 | 0.529 |
0.3 | 4.354 | 4.260 | 4.004 | 0.881 | 0.879 | 0.905 | 0.509 |
0.5 | 4.413 | 4.388 | 4.116 | 0.905 | 0.872 | 0.908 | 0.441 |
Inspired by LLM planning literature such as ReAct Yao et al. (2022) and ExpeL Zhao et al. (2023), we incorporate an experience pool into our algorithm to enhance molecular optimization. The experience pool consists of two key components: (1) valuable insights and important information kept from past experience, and (2) a summary is constructed from both high-performing and suboptimal molecules generated during the optimization process. These are achieved by summarizing information from the top 10 and bottom 10 molecules in each iteration. The worst 10 molecules are selected using a sliding window approach with a stride of 10. Specifically, if in the previous iteration, the worst molecules were extracted from the bottom 10, the next iteration extracts molecules ranked from the bottom 20 to the bottom 10. This mechanism ensures that the experience pool continuously evolves, integrating knowledge from both current and past iterations. However, incorporating experience into every generation step carries a high risk of convergence to local optima, as it overly exploits the currently discovered search directions. To address this concern, we conduct additional experiments by varying the probability of incorporating experience, as shown in Table 6 and determine to use 0.5. A detailed analysis comparing the initial and final experience in the experience pool is provided in Appendix 7.1. We also visualize the convergence behavior of the top-10 and top-100 molecules generated by baseline models in Figure 2. MOLLM exhibits highly stable training curves, consistently outperforming all other models with a substantial margin over the second-best. Notably, even without the experience pool, MOLLM surpasses MOLLEO, highlighting the effectiveness of its other components.
6 Conclusion
In this work, we introduce MOLLM, a novel framework that integrates MOO, GA, and LLMs with in-context learning and prompt engineering for molecular design. MOLLM requires no additional training, relying exclusively on LLMs as genetic operators, and achieves SOTA performance in multi-objective molecular optimization and PMO benchmark. Through rigorous framework design, empirical evaluations and analysis, we demonstrate its effectiveness and efficiency. MOLLM significantly reduces computational costs while outperforming other LLM-based approaches and other SOTA methods. This efficiency is particularly advantageous for practical applications, where molecular property evaluations often involve costly biological and pharmaceutical testing, and LLM inference imposes a high computational overhead. Our results show that MOLLM maintains robust performance across various objective settings and remains superior when optimizing many objectives, including less predictable properties such as BBBP. Furthermore, MOLLM is adaptable to various open-source and closed-source LLMs and can be readily extended to other domains with minor modifications to the our code. The source code will be released as open source upon acceptance of the paper to facilitate reproducibility and further research. Due to the limit of pages, we will also provide the ADMET analysis of MOLLM upon acceptance of the paper. However, one limitation of our work is that relying solely on a single LLM to generate new molecules may increase the risk of convergence to local optima. Future research may focus on enhancing the diversity of search directions and improving the exploration efficiency of MOLLM.